# Faster Online Computation of String Net Frequency

### v1

@@ Abstract
An occurrence of a repeated substring $u$ in a string $S$ is called a net occurrence if extending the occurrence to the left or to the right decreases the number of occurrences to 1. The net frequency (NF) of a repeated substring $u$ in a string $S$ is the number of net occurrences of $u$ in $S$. Very recently, Guo et al. [SPIRE 2024] proposed an online $O(n\log\sigma)$-time algorithm that maintains a data structure of $O(n)$ space which answers Single-NF queries in $O(m\log\sigma+\sigma^{2})$ time and reports all answers of the All-NF problem in $O(n\sigma^{2})$ time. Here, $n$ is the length of the input string $S$, $m$ is the query pattern length, and $\sigma$ is the alphabet size. The $\sigma^{2}$ term is a major drawback of their method since computing string net frequencies is originally motivated for Chinese language processing where $\sigma$ can be thousands large. This paper presents an improved online $O(n\log\sigma)$-time algorithm, which answers Single-NF queries in $O(m\log\sigma)$ time and reports all answers to the All-NF problem in output-optimal $O(|\mathsf{NF^{+}}(S)|)$ time, where $\mathsf{NF^{+}}(S)$ is the set of substrings of $S$ paired with their positive NF values. We note that $|\mathsf{NF^{+}}(S)|=O(n)$ always holds. In contract to Guo et al.’s algorithm that is based on Ukkonen’s suffix tree construction, our algorithm is based on Weiner’s suffix tree construction.
@@ 1 Introduction
In this paper, we present an improved algorithm for the Single-NF and All-NF problems in the online setting (see also Table 1). We present an online algorithm that uses $O(n)$ space and maintains a list of substrings with positive NF values in $O(\log\sigma)$ amortized time per character, for general ordered alphabets of size $\sigma$. Since the positive NF values can be stored in the corresponding nodes in the suffix tree with $O(n)$ space, our data structure answers Single-NF queries in $O(m\log\sigma)$ time each. Also, our algorithm maintains a data structure that can report all answers to the All-NF problem in output optimal time, independently of the alphabet size $\sigma$.
@@ 2.2 String Net Frequency and Maximal Repeats
A string $u$ with $\mathsf{occ}_{S}(u)\geq 2$ is called a repeat in $S$. For a repeat $u\in\mathsf{Substr}(S)$ of a string $S$, let
$\Phi_{S}(u)=\{(a,b)\in\Sigma\times\Sigma\mid\mathsf{occ}_{S}(aub)=\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1\}.$
The net frequency (NF) of $u$ in $S$, denoted $\phi_{S}(u),$ is the number of ordered pairs $(a,b)\in\Sigma\times\Sigma$ of characters such that $\mathsf{occ}_{S}(aub)=\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1$, i.e., $\phi_{S}(u)=|\Phi_{S}(u)|$.
A repeat $u$ in a string $S$ is called a left-maximal repeat in $S$ if $u$ is a prefix of $S$, or there are at least two distinct characters $a,a^{\prime}$ that immediately precede $u$ in $S$ (i.e. $au,a^{\prime}b\in\mathsf{Substr}(S)$). Similarly, a repeat $u$ in $S$ is called a right-maximal repeat in $S$ if $u$ is a suffix of $S$, or there are at least two distinct characters $b,b^{\prime}$ that immediately follow $u$ in $S$ (i.e. $ub,ub^{\prime}\in\mathsf{Substr}(S)$). Let $\mathsf{LM}(S)$ and $\mathsf{RM}(S)$ denote the sets of left-maximal and right-maximal repeats in $S$, respectively.
For a left-maximal repeat $u$ in $S$, a character $a\in\Sigma$ such that $\mathsf{occ}_{S}(au)\geq 1$ is called a left-extension of $u$ in $S$. Similarly, for a right-maximal repeat $u$ in $S$, a character $b\in\Sigma$ such that $\mathsf{occ}_{S}(ub)\geq 1$ is called a right-extension of $u$ in $S$.
A repeat $u$ is called a maximal repeat in $S$ if $u\in\mathsf{LM}(S)\cap\mathsf{RM}(S)$. Let $\mathsf{M}(S)=\mathsf{LM}(S)\cap\mathsf{RM}(S)$ denote the set of maximal repeats in $S$.
Guo et al. [8] showed that if $\phi_{S}(u)\geq 1$, then $u\in\mathsf{RM}(S)$. Below we present a stronger statement.
@@
Since $\phi_{S}(u)\geq 1$, there exists a pair $(a,b)$ of characters such that $\mathsf{occ}_{S}(au)=\mathsf{occ}_{S}(ub)=1$. Since $\mathsf{occ}_{S}(u)\geq 2$, there is an occurrence $u=S[i..i+|u|-1]$ of $u$ such that $S[i-1]\neq a$ or $i=1$. Similarly, there is an occurrence $u=S[j-|u|+1..j]$ of $u$ such that $S[j+1]\neq b$ or $j=|S|$. Thus $u\in\mathsf{LM}(S)\cap\mathsf{RM}(S)=\mathsf{M}(S)$. ∎
@@
$\mathsf{NF^{+}}(S)=\{(u,\phi_{S}(u))\mid u\in\mathsf{Substr}(S),\phi_{S}(u)\geq 1\}$
denote the set of pairs of string $u$ and its NF that is positive. It follows from Lemma 1 that if $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$, then $u\in\mathsf{M}(S)$. Since $|\mathsf{M}(S)|\leq n-1$ for any string of length $n$ (c.f. [3]), it is immediate that $|\mathsf{NF^{+}}(S)|\leq n-1$.
@@ 2.3 Online Computation of String Net Frequency
We will design an online algorithm for computing substrings that have positive values of net frequencies, in which new characters are prepended to the input string. Namely, our algorithm works in a right-to-left online manner. Note that $\phi_{S}(u)$ is closed under reversal, i.e., $\phi_{S}(u)=\phi_{S^{R}}(u^{R})$. Hence, by reversing the input string, our algorithm can be used in the left-to-right online setting as well.
We consider the following problems of computing string net frequencies online.
@@
Let $S$ be a string of length $n$ for which $\mathsf{NF^{+}}(S)$ has been computed. Given a character $a\in\Sigma$, compute $\mathsf{NF^{+}}(aS)\Delta\mathsf{NF^{+}}(S)=\left(\mathsf{NF^{+}}(aS)\setminus\mathsf{NF^{+}}(S)\right)\cup\left(\mathsf{NF^{+}}(S)\setminus\mathsf{NF^{+}}(aS)\right)$.
Problem 1 requires one to perform the following per new character:
  1. (i)
    Update the NFs of some existing substrings.
  2. (i)
    Add substrings whose NFs become positive;
  3. (iii)
    Remove substrings whose NFs become 0;
@@ 3 Algorithm
Let $S$ be the string. As in the previous work [9], a node $u$ in $\mathsf{STree}(S)$ stores the NF value $\phi_{S}(u)$ iff $\phi_{S}(u)\geq 1$.
@@ 3.1 Key Lemma
Let $u$ be a repeat in a string $S$. Let $(a,b)\in\Phi_{S}(u)$. Then, for any character $c\neq b$, $(a,c)\notin\Phi_{S}(u)$. Similarly, for any character $d\neq a$, $(d,b)\notin\Phi_{S}(u)$.
@@
For a contradiction assume that $(a,c)\in\Phi_{S}(u)$. Then, $\mathsf{occ}_{S}(aub)=1$ and $\mathsf{occ}_{S}(auc)=1$. Since $b\neq c$, $\mathsf{occ}_{S}(au)\geq 2$. However, since $(a,b)\in\Phi_{S}(u)$, we have $\mathsf{occ}_{S}(au)=1$, a contradiction. The case for $(d,b)$ is symmetric. ∎
Lemma 3 implies that for any character $a\in\Sigma$ that is a left-extension of a (maximal) repeat $u$ in $S$, there is at most one character $b\in\Sigma$ that contributes to the positive NF of $u$ in $S$. Thus the following corollary is immediate:
@@
For any repeat $u$ in a string $S$, $\phi_{S}(u)\leq\min\{|\Sigma^{L}_{u}|,|\Sigma^{R}_{u}|\}$, where $\Sigma^{L}_{u}$ and $\Sigma^{R}_{u}$ are the sets of left-extensions and right-extensions of $u$ in $S$, respectively.
Recall that the right-extensions of (right) maximal repeats are the first characters of the edge labels of $\mathsf{STree}(S)$. Since any string of length $n$ can contain at most $n-2$ non-empty maximal repeats [3], the next lemma follows from Lemma 1 and Corollary 1.
@@
For any string $S$ of length $n$, $\sum_{u\in\mathsf{Substr}(S)}\phi_{S}(u)\leq n-2$.
This is a slight improvement over the upper bound $\sum_{u\in\mathsf{Substr}(S)}\phi_{S}(u)\leq n$ given by Guo et al. [8].

### v2

# Faster and Simpler Online Computation of String Net Frequency

@@ Abstract
An occurrence of a repeated substring $u$ in a string $S$ is called a net occurrence if extending the occurrence to the left or to the right decreases the number of occurrences to 1. The net frequency (NF) of a repeated substring $u$ in a string $S$ is the number of net occurrences of $u$ in $S$. Very recently, Guo et al. [SPIRE 2024] proposed an online $O(n\log\sigma)$-time algorithm that maintains a data structure of $O(n)$ space which answers Single-NF queries in $O(m\log\sigma+\sigma^{2})$ time and reports all answers of the All-NF problem in $O(n\sigma^{2})$ time. Here, $n$ is the length of the input string $S$, $m$ is the query pattern length, and $\sigma$ is the alphabet size. The $\sigma^{2}$ term is a major drawback of their method since computing string net frequencies is originally motivated for Chinese language text processing where $\sigma$ can be thousands large. This paper presents an improved online $O(n\log\sigma)$-time algorithm, which answers Single-NF queries in $O(m\log\sigma)$ time and reports all answers to the All-NF problem in output-optimal $O(|\mathsf{NF^{+}}(S)|)$ time, where $\mathsf{NF^{+}}(S)$ is the set of substrings of $S$ paired with their positive NF values. We note that $|\mathsf{NF^{+}}(S)|=O(n)$ always holds. In contract to Guo et al.’s algorithm that is based on Ukkonen’s suffix tree construction, our algorithm is based on Weiner’s suffix tree construction.
@@ 1 Introduction
In this paper, we resolve both of the aforementioned issues by presenting an improved algorithm for the Single-NF and All-NF problems in the online setting (see also Table 1). We present an online algorithm that uses $O(n)$ space and maintains a list of repeats with positive NF values in $O(\log\sigma)$ amortized time per character, for general ordered alphabets of size $\sigma$. Since the positive NF values can be stored in the corresponding nodes in the suffix tree with $O(n)$ space, our data structure answers Single-NF queries in $O(m\log\sigma)$ time each. Also, our algorithm maintains a data structure that can report all answers to the All-NF problem in output optimal time, independently of the alphabet size $\sigma$.
@@
Very recently, Mieno and Inenaga [14] proposed another online NF algorithm that is based on Ukkonen’s left-to-right suffix tree construction [17]. While the MI algorithm of [14] requires some non-trivial tweaks for maintaining the secondary active point on top of the usual active point over the online suffix tree, no such additional mechanism is required in our Weiner-based method.
@@ 2.2 String Net Frequency and Maximal Repeats
A string $w$ with $\mathsf{occ}_{S}(w)\geq 2$ is called a repeat in $S$. For a repeat $w\in\mathsf{Substr}(S)$ of a string $S$, let
$\Phi_{S}(w)=\{(\alpha,\beta)\in\Sigma\times\Sigma\mid\mathsf{occ}_{S}(\alpha w\beta)=\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=1\}.$ (1)
The net frequency (NF) of $w$ in $S$, denoted $\phi_{S}(w),$ is the number of ordered pairs $(\alpha,\beta)\in\Sigma\times\Sigma$ of characters such that $\mathsf{occ}_{S}(\alpha w\beta)=\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=1$, i.e.,
$\phi_{S}(w)=|\Phi_{S}(w)|.$ (2)
A repeat $w$ in a string $S$ is called a left-maximal repeat in $S$ if $w$ is a prefix of $S$, or there are at least two distinct characters $\alpha,\alpha^{\prime}$ that immediately precede $w$ in $S$ (i.e. $\alpha w,\alpha^{\prime}w\in\mathsf{Substr}(S)$). Similarly, a repeat $w$ in $S$ is called a right-maximal repeat in $S$ if $w$ is a suffix of $S$, or there are at least two distinct characters $\beta,\beta^{\prime}$ that immediately follow $w$ in $S$ (i.e. $w\beta,w\beta^{\prime}\in\mathsf{Substr}(S)$). Let $\mathsf{LM}(S)$ and $\mathsf{RM}(S)$ denote the sets of left-maximal and right-maximal repeats in $S$, respectively.
For a left-maximal repeat $u$ in $S$, a character $\alpha\in\Sigma$ such that $\mathsf{occ}_{S}(\alpha w)\geq 1$ is called a left-extension of $w$ in $S$. Similarly, for a right-maximal repeat $w$ in $S$, a character $\beta\in\Sigma$ such that $\mathsf{occ}_{S}(w\beta)\geq 1$ is called a right-extension of $w$ in $S$.
A repeat $u$ is called a maximal repeat in $S$ if $w\in\mathsf{LM}(S)\cap\mathsf{RM}(S)$. Let $\mathsf{M}(S)=\mathsf{LM}(S)\cap\mathsf{RM}(S)$ denote the set of maximal repeats in $S$.
Guo et al. [8] showed that if $\phi_{S}(w)\geq 1$, then $w\in\mathsf{RM}(S)$. Below we present a stronger statement.
@@
Since $\phi_{S}(w)\geq 1$, there exists a pair $(\alpha,\beta)$ of characters such that $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=1$. Since $\mathsf{occ}_{S}(w)\geq 2$, there is an occurrence $w=S[i..i+|w|-1]$ of $u$ such that $S[i-1]\neq\alpha$ or $i=1$. Similarly, there is an occurrence $w=S[j-|w|+1..j]$ of $w$ such that $S[j+1]\neq\beta$ or $j=|S|$. Thus $w\in\mathsf{LM}(S)\cap\mathsf{RM}(S)=\mathsf{M}(S)$. ∎
@@
$\mathsf{NF^{+}}(S)=\{(w,\phi_{S}(w))\mid w\in\mathsf{M}(S),\phi_{S}(w)\geq 1\}$
denote the set of pairs of string $w$ and its NF that is positive in $S$. It follows from Lemma 1 that if $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$, then $w\in\mathsf{M}(S)$. Since $|\mathsf{M}(S)|\leq n-1$ for any string of length $n$ (c.f. [3]), it is immediate that $|\mathsf{NF^{+}}(S)|\leq n-1$.
@@ 2.3 Online Computation of String Net Frequency
We will design an online algorithm for computing substrings that have positive values of net frequencies, in which new characters are prepended to the input string. Namely, our algorithm works in a right-to-left online manner. Note that $\phi_{S}(w)$ is closed under reversal, i.e., $\phi_{S}(w)=\phi_{S^{R}}(w^{R})$. Hence, by reversing the input string, our algorithm can be used in the left-to-right online setting as well.
We consider the following problem of maintaining the list $\mathcal{L}$ that stores $\mathsf{NF^{+}}(S)$ for an online string $S$:
@@
Let $S$ be a string of length $n$ for which $\mathsf{NF^{+}}(S)$ has been computed. Given a character $a\in\Sigma$, update $\mathcal{L}$ by computing $\mathsf{NF^{+}}(aS)\Delta\mathsf{NF^{+}}(S)=\left(\mathsf{NF^{+}}(aS)\setminus\mathsf{NF^{+}}(S)\right)\cup\left(\mathsf{NF^{+}}(S)\setminus\mathsf{NF^{+}}(aS)\right)$.
Problem 1 requires one to consider the following cases each time a new character $a\in\Sigma$ is prepended to the current string $S$:
Case (i)
Increase the NFs of some existing maximal repeats: For each $w\in\mathsf{M}(aS)$ such that $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)>\phi_{S}(w)\geq 1$, replace $(w,\phi_{S}(w))$ with $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$.
Case (ii)
Add maximal repeats whose NFs become positive: For each $w\in\mathsf{M}(aS)$ such that $(w,\phi_{S}(w))\notin\mathsf{NF^{+}}(S)$, $\phi_{S}(w)=0$, and $\phi_{aS}(w)\geq 1$, add $(w,\phi_{aS}(w))$ to the list $\mathcal{L}$.
Case (iii)
Decrease the NFs of some existing maximal repeats: For each $w\in\mathsf{M}(aS)$ such that $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)<\phi_{S}(w)$, replace $(w,\phi_{S}(w))$ with $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$ if $\phi_{aS}(w)\geq 1$, and remove $(w,\phi_{S}(w))$ from $\mathcal{L}$ if $\phi_{aS}(w)=0$.
Example 1.
Consider string $S=\mathtt{ababbababcababbb\$}$ of which the set of maximal repeats is $\mathsf{M}(S)=\{\mathtt{b},\mathtt{ab},\mathtt{bb},\mathtt{bab},\mathtt{abab},\mathtt{ababb}\}$. We have
  • $\Phi_{S}(\mathtt{bb})=\{(\underline{\mathtt{b}},\underline{\mathtt{\$}})\}$ and thus $\phi_{S}(\mathtt{bb})=1$, because $\mathsf{occ}_{S}(\mathtt{bb})=3$ and $\mathsf{occ}_{S}(\mathtt{\underline{b}bb})=\mathsf{occ}_{S}(\mathtt{bb\underline{\$}})=\mathsf{occ}_{S}(\mathtt{\underline{b}bb\underline{\$}})=1$;
  • $\Phi_{S}(\mathtt{bab})=\{(\mathtt{\underline{b}},\mathtt{\underline{a}})\}$ and thus $\phi_{S}(\mathtt{bab})=1$, because $\mathsf{occ}_{S}(\mathtt{bab})=4$ and $\mathsf{occ}_{S}(\mathtt{\underline{b}bab})=\mathsf{occ}_{S}(\mathtt{bab\underline{a}})=\mathsf{occ}_{S}(\mathtt{\underline{b}bab\underline{a}})=1$;
  • $\Phi_{S}(\mathtt{abab})=\{(\mathtt{\underline{b}},\mathtt{\underline{c}})\}$ and thus $\phi_{S}(\mathtt{abab})=1$, because $\mathsf{occ}_{S}(\mathtt{abab})=3$ and $\mathsf{occ}_{S}(\mathtt{\underline{b}abab})=\mathsf{occ}_{S}(\mathtt{abab\underline{c}})=\mathsf{occ}_{S}(\mathtt{\underline{b}abab\underline{c}})=1$;
  • $\Phi_{S}(\mathtt{ababb})=\{(\mathtt{\underline{c}},\mathtt{\underline{b}})\}$ and thus $\phi_{S}(\mathtt{ababb})=1$, because $\mathsf{occ}_{S}(\mathtt{ababb})=2$ and $\mathsf{occ}_{S}(\mathtt{\underline{c}ababb})=\mathsf{occ}_{S}(\mathtt{ababb\underline{b}})=\mathsf{occ}_{S}(\mathtt{\underline{c}ababb\underline{b}})=1$.
Since the other maximal repeats in $S$ do not have positive net frequencies, the list $\mathcal{L}$ stores $\mathsf{NF^{+}}(S)=\{(\mathtt{bb},1),(\mathtt{bab},1),(\mathtt{abab},1),(\mathtt{ababb},1)\}$.
Consider string $\mathtt{b}S=\mathtt{bababbababcababbb\$}$ of which the set of maximal repeats is $\mathsf{M}(\mathtt{b}S)=\{\mathtt{b},\mathtt{ab},\mathtt{bb},\mathtt{bab},\mathtt{abab},\mathtt{ababb},\mathtt{babab}\}$. For the new maximal repeat $\mathtt{babab}\in\mathsf{M}(\mathtt{b}S)\setminus\mathsf{M}(S)$ we have
  • $\Phi_{\mathtt{b}S}(\mathtt{babab})=\{(\mathtt{\underline{b},\mathtt{\underline{c}}})\}$ and thus $\phi_{\mathtt{b}S}(\mathtt{babab})=1$, because $\mathsf{occ}_{\mathtt{b}S}(\mathtt{babab})=2$ and $\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}babab})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{babab\underline{c}})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}babab\underline{c}})=1$. This is in Case (ii).
For the existing maximal repeats in $\mathsf{M}(S)\cap\mathsf{M}(\mathtt{b}S)$, we have
  • $\Phi_{\mathtt{b}S}(\mathtt{ababb})=\{(\mathtt{\underline{b}},\mathtt{\underline{a}}),(\mathtt{b},\mathtt{c})\}$ and thus $\phi_{\mathtt{b}S}(\mathtt{abab})=2$, because $\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababb})=2$, $\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}ababb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababb\underline{a}})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}ababb\underline{a}})=1$, and $\mathsf{occ}_{\mathtt{b}S}(\mathtt{cababb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababbb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{cababbb})=1$. This is in Case (i);
  • $\Phi_{\mathtt{b}S}(\mathtt{abab})=\emptyset$ and thus $\phi_{\mathtt{b}S}(\mathtt{abab})=0$, because $\mathsf{occ}_{\mathtt{b}S}(\mathtt{abab})=3$ and $\mathsf{occ}_{\mathtt{b}S}(\mathtt{abab\underline{c}})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}abab\underline{c}})=1$ but $\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}abab})=2$. This is in Case (iii);
  • $\Phi_{\mathtt{b}S}(\mathtt{bab})=\emptyset$ and thus $\phi_{\mathtt{b}S}(\mathtt{bab})=0$, because $\mathsf{occ}_{\mathtt{b}S}(\mathtt{bab})=4$ and $\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}bab})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}bab\underline{a}})=1$ but $\mathsf{occ}_{\mathtt{b}S}(\mathtt{bab\underline{a}})=2$. This is in Case (iii).
Since there are no other changes, the elements in the list $\mathcal{L}$ are updated to $\mathsf{NF^{+}}(\mathtt{b}S)=\{(\mathtt{bb},1),(\mathtt{ababb},2),(\mathtt{babab},1)\}$.
@@ 2.4 Suffix Trees and Weiner Links
Throughout the description of our algorithm in Section 3, we will assume that the input string $S$ terminates with a unique end-marker \$currency-dollar\\$\$. We note that this assumption only gives little impact in our NF computation. By our key lemma (Lemma 3) from the next section, we have that there can exist at most one character $\alpha$ such that $(\alpha,\$)\in\Phi_{S}(w)$ for a repeat $w$ in $S$. Moreover, there can exist at most one repeat $w$ that can have \$currency-dollar\\$\$ in the second element of the pair $(\alpha,\$)\in\Phi_{S}(w)$: Let $w$ be the longest repeated suffix of $S[1..n-1]$ where $n=|S|$, and let $\alpha w$ be the shortest non-repeated suffix of $S[1..n-1]$ where $\alpha\in\Sigma$. Since $S[n]=\$$, we have $\mathsf{occ}_{S[1..n-1]}(w)=\mathsf{occ}_{S}(w)>1$ and $\mathsf{occ}_{S[1..n-1]}(\alpha w)=\mathsf{occ}_{S}(\alpha w)=1$. Since $\mathsf{occ}_{S}(w\$)=\mathsf{occ}_{S}(\alpha w\$)=1$, $(\alpha,\$)\in\Phi_{S}(w)$ holds. Note that no other suffixes of $S[1..n-1]$ can satisfy these properties.
@@ 3 Algorithm
Let $S$ be the input string. As in the previous work [9], a node $w$ in $\mathsf{STree}(S)$ stores the NF value $\phi_{S}(w)$ iff $\phi_{S}(w)\geq 1$.
@@ 3.1 Key Lemma
Let $w$ be a repeat in a string $S$. Let $(\alpha,\beta)\in\Phi_{S}(w)$. Then, for any character $\beta^{\prime}\neq\beta$, $(\alpha,\beta^{\prime})\notin\Phi_{S}(w)$. Similarly, for any character $\alpha^{\prime}\neq\alpha$, $(\alpha^{\prime},\beta)\notin\Phi_{S}(w)$.
@@
For a contradiction assume that $(\alpha,\beta^{\prime})\in\Phi_{S}(w)$. Then, $\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{S}(\alpha w\beta^{\prime})=1$. Since $\beta\neq\beta^{\prime}$, $\mathsf{occ}_{S}(\alpha w)\geq 2$. However, since $(\alpha,\beta)\in\Phi_{S}(w)$, we have $\mathsf{occ}_{S}(\alpha w)=1$, a contradiction. The case for $(\alpha^{\prime},\beta)$ is symmetric. ∎
Lemma 3 implies that for any character $\alpha\in\Sigma$ that is a left-extension of a (maximal) repeat $w$ in $S$, there is at most one character $\beta\in\Sigma$ that contributes to the positive NF of $w$ in $S$. Thus the following corollary is immediate:
@@
For any repeat $w$ in a string $S$, $\phi_{S}(w)\leq\min\{|\Sigma^{L}_{w}|,|\Sigma^{R}_{w}|\}$, where $\Sigma^{L}_{w}$ and $\Sigma^{R}_{w}$ are the sets of left-extensions and right-extensions of $u$ in $S$, respectively.
Since the total number of right-extensions of the maximal repeats is most $22n-$ 2 for any string of length $n\geq 2$ (c.f. [3]), the next lemma immediately follows from Lemma 1 and Corollary 1.
@@
For any string $S$ of length $n$, $\sum_{w\in\mathsf{Substr}(S)}\phi_{S}(u)\leq 2n-2$.
This bound is weaker than the bound $\sum_{w\in\mathsf{Substr}(S)}\phi_{S}(w)\leq n$ given by Guo et al. [8], but it is sufficient for our purpose. Namely, if needed, one can explicitly store the character pairs in $\Phi_{S}(w)$ in $O(n)$ total space instead of the counts $\phi_{S}(w)$. Our algorithm to follow can easily be modified to store the character pairs in $\Phi_{S}(w)$ as well, in the same complexity.